A fair six-sided dice is thrown and the scores are noted.

Event X: The total of the two scores is 4.

Even Y: The first score is 2 or 5.                                       

b)             Find the probability of X and Y.                                                                        


1
Expert's answer
2021-06-28T16:26:28-0400

A dice is thrown two times. So, total outcomes = "6^2=36."

Event X: The total of the two scores is 4


"(1,3), (2,2), (3,1)"

"P(X)=\dfrac{3}{36}=\dfrac{1}{12}"

Event Y: The first score is 2 or 5


"(2,1), (2,2), (2,3),(2,4), (2,5), (2,6),"

"(5,1), (5,2), (5,3),(5,4), (5,5), (5,6),"

"P(Y)=\dfrac{12}{36}=\dfrac{1}{3}"